OpenStudy (anonymous):

use a graphing utility to find the sum of each geometric sequence. (1/4) + (2/4) + (2^2/4) + (2^3/4) + ... + (2^14/4)

terenzreignz (terenzreignz):

Please tell me you didn't try to do this manually :D

OpenStudy (anonymous):

lol im very close to....

terenzreignz (terenzreignz):

It's a geometric series, and if written in the convenient sigma notation, it'd look something like this... \[\huge \sum_{n=1}^{15}\frac{2^{n-1}}4\]

OpenStudy (anonymous):

now the answer is in the back of the book and it shows me everything to put into my calculator and gives me the answer but it doesn't tell me where the buttons are im supposed to be pressing

OpenStudy (anonymous):

I wrote that formula down next to the problem

terenzreignz (terenzreignz):

Oh... can't help you with that, I have no idea what calculator you're using... There is a formula for geometric series...though...

OpenStudy (anonymous):

the issue im having is it says to solve by using a calculator and I don't know how to put it into that

OpenStudy (anonymous):

im using a TI84 plus

terenzreignz (terenzreignz):

Yikes... better inquire as to who has such a calculator, then :D Good luck :)

OpenStudy (anonymous):

8191.75

OpenStudy (anonymous):

the answer is 8191.75 I have the answer and what is put into the calculator to get it I just don't know what to put down in the calculator

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

now what do you put into your calculator I have a TI 84

OpenStudy (anonymous):

the sum of GP series is \[a(r ^{n}-1)/(r-1)\] here a is 2, r=2, n=15... then divide it by 8 as the equation is \[\sum_{n=1}^{15} 2^{n}/8\]

terenzreignz (terenzreignz):

whoa there... a is not 2... a is the first term in this progression...

OpenStudy (anonymous):

i dont have a have a graph calculator... @terenzreignz , in this equation, 'a' a.k.a 'the first term' is 2

terenzreignz (terenzreignz):

No... it's 1/4

OpenStudy (anonymous):

okay wait why is the number of the summation sign 15 and not 14

terenzreignz (terenzreignz):

Because there would be 15 terms, @d2thej

OpenStudy (anonymous):

if the number stops at 14 wouldn't there only be 14?

OpenStudy (anonymous):

do you just add one to that number just for the heck of it?

OpenStudy (anonymous):

simplifiy the equation you have written in the previous answer.......2^(n-1)/4......make it (2^n)/8

terenzreignz (terenzreignz):

It does stop at 14 (the exponent of 2) however, it STARTED at 0. See, the first term is 1/4 OR \[\huge \frac{2^0}{4}\] So there are, in fact, 15 terms...

OpenStudy (anonymous):

hmm I guess ill jus take your word for it, im not seeing it

OpenStudy (anonymous):

@jiraya what do you mean what am I simplifying

terenzreignz (terenzreignz):

Okay... visualize :D Let's stop it at 3, for instance... \[\huge \frac{1}{4}\quad\frac{2}{4}\quad\frac{2^2}{4}\quad\frac{2^3}{4}\] See, when we stop at 3, we, in fact, have 4 terms... similar logic, basically

OpenStudy (anonymous):

ahhhhhh okay

OpenStudy (anonymous):

so its like you count how many are raised to a power and the one that isn't I guess?

terenzreignz (terenzreignz):

Of course :) We're counting TERMS and that's that, no matter what they look like. But if you really must have uniformity... \[\huge \frac{2^0}{4}\quad\frac{2^1}{4}\quad\frac{2^2}{4}\quad\frac{2^3}{4}...\quad\frac{2^{14}}{4}\] And now it's clear to you that there are 15 terms? Because we stopped at 14, but we STARTED at 0 (not 1)

OpenStudy (anonymous):

yes! I get it now lol thanks, so what buttons am I pushing exactly to put it in my calculator lol

terenzreignz (terenzreignz):

That... you're gonna have to find another way to do that, I don't have your calculator, sorry bro :D

OpenStudy (anonymous):

nvm I seen what @jiraya did

OpenStudy (anonymous):

lol thanks though

OpenStudy (anonymous):

okay wait now,

OpenStudy (anonymous):

how would I know to put 8 on the bottom instead of 4

terenzreignz (terenzreignz):

I guess it's just to make the summation look nicer... hang on...

OpenStudy (anonymous):

okay I got more questions if you don't mind lol

terenzreignz (terenzreignz):

I have to go in a bit... maybe just one :)

OpenStudy (anonymous):

okay same instructions except they give me this: \[\sum_{n=1}^{15} (2/3)^n\]

terenzreignz (terenzreignz):

Even better (that is, if you're trying to find the sum) You can just plug in directly :D

terenzreignz (terenzreignz):

Obviously your common ratio r = 2/3 and your a (first term) would be (2/3)^1 = 2/3 and just plug in

OpenStudy (anonymous):

it doesn't come out to what the answer is in back of the book lol

terenzreignz (terenzreignz):

What formula did you use?

OpenStudy (anonymous):

the summation one

OpenStudy (anonymous):

ohhh wait nvm I got it

terenzreignz (terenzreignz):

This formula... \[\Large \frac{a(1-r^n)}{1-r}\]

OpenStudy (anonymous):

it depended on how I put it into the calculator

OpenStudy (anonymous):

i don't know if you've ever used one of these before but the 2/3 had to be in its separate parenthesis other than what the equation gives you

OpenStudy (anonymous):

thanks for your help if you need to go that's cool lol can i add you as a friend on here?

terenzreignz (terenzreignz):

Seems you're way ahead of me :) I really need to go now, though, try to find someone who knows your calculator inside and out :D I'll be seeing you :) ------------------------------------------------------ Terence out